# Self-Data Distillation for Recovering Quality in Pruned Large Language Models

### v2

@@ 6 Conclusion
In conclusion, we introduced self-data distilled fine-tuning as an effective method to counteract quality degradation in pruned Llama3.1-8B Instruct models, addressing catastrophic forgetting while preserving alignment with the model’s original distribution. Our approach consistently outperforms standard supervised fine-tuning, demonstrating superior accuracy recovery post-pruning across various downstream tasks on the HuggingFace OpenLLM Leaderboard v1. Additionally, model merging via SLERP further enhances recovery, achieving significant quality retention. We also show that our method scales with dataset size, where larger self-distilled datasets lead to improved quality recovery. These findings highlight self-data distilled fine-tuning as a critical tool for maintaining high model quality post-pruning, offering an efficient solution for large-scale model compression. Future work may involve integrating self-data distilled fine-tuning with complementary model compression techniques such as sparsity, quantization or teacher distillation, potentially yielding greater efficiency without sacrificing model quality. Furthermore, adopting fine-tuning strategies that leverage dynamically generated datasets or incorporate multi-modal inputs could enhance the retention of critical knowledge, broadening the scope of self-data distillation to a wider array of tasks. Extending these methodologies to next-generation LLM architectures presents a promising avenue for unlocking additional computational efficiency and model robustness.
7 Acknowledgements
We would like to express our gratitude to Valavan Manohararajah, Mike Lasby, Gokul Ramakrishnan, Alex Tsaptsinos, and Eric Sather for their valuable discussions, as well as their insightful comments and feedback, which greatly enhanced the quality of our manuscript.
@@ A.1 Baseline Model
We used Llama3.1-8B Instruct [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct(Dubey et al., 2024) as the baseline model for all experiments. This model comprises a total of 32 decoder layers, pretrained on a diverse array of instruction-following datasets. This model was chosen for its strong generalization performance across a wide range of natural language processing (NLP) tasks, making it an ideal candidate for studying the impact of structured pruning and fine-tuning. The 8B model size strikes a balance between computational efficiency and model quality, providing a robust foundation for the experiments in this study. Hence, served as the starting point for our structured layer pruning ablations and experiments in Sections 3 and 4, respectively.
@@ A.2 Structured Layer Pruning
A.3 Fine-tuning Datasets
@@
  • Alpaca Cleaned [^2] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023) The Alpaca Cleaned dataset is a cleaned version of the original Stanford Alpaca dataset, containing 50,000 instruction-following examples. It addresses several issues present in the original release, such as hallucinations, incorrect instructions, and output inconsistencies. This dataset provides high-quality general instruction-following tasks, spanning text generation, summarization, reasoning, and more. The cleaned version offers improved consistency and accuracy, making it ideal for fine-tuning large models in real-world instruction-following tasks.
  • OpenMathInstruct (Toshniwal et al., 2024)T The OpenMathInstruct-1 dataset is specifically designed for fine-tuning language models on mathematical instruction tasks. It contains 1.8 million problem-solution pairs, generated using Mixtral-8x7B (Jiang et al., 2024). The problem sets are drawn from well-established mathematical benchmarks, including the GSM8K and MATH (Hendrycks et al., 2021b) datasets, ensuring a diverse and challenging range of mathematical reasoning tasks. Solutions are generated synthetically by allowing the Mixtral model to leverage a combination of natural language reasoning and executable Python code, which allows for both symbolic computation and procedural solutions. This combination of text and code execution makes the dataset particularly suited for training models to handle complex reasoning, problem-solving, and algebraic tasks.
A.3.1 Data Sampling and Experimental Consistency
@@
A.4 Fine-tuning Pruned Models
@@
Through this grid search, the optimal configuration was identified as a rank size = 8, epochs = 5, a batch size = 64, and learning rate = 1$\times$10-4. These hyperparameters were used consistently across all fine-tuning experiments (i.e., both standard supervised fine-tuning and self-data distilled fine-tuning) in this study to ensure a fair comparison of the models and their quality post-pruning. We conduct our model training using LLaMA-Factory v0.8.3333https://github.com/hiyouga/LLaMA-Factory/releases/tag/v0.8.3, a versatile framework designed for large-scale language model training and fine-tuning. This version offers extensive support for efficient parallelism, optimized memory usage, and integration with popular datasets, making it ideal for large model fine-tuning tasks such as those performed in this study.
A.4.1 Computational Resources
@@ Appendix C Experimental Setup for Understanding Catastrophic Forgetting
To understand the impact of distribution shift on catastrophic forgetting, we conducted experiments using the baseline model (i.e., Llama3.1-8B Instruct) and its pruned variants fine-tuned with both supervised fine-tuning (SFT) and self-data distilled fine-tuning (Self-Data FT). Specifically, we pruned 6 decoder layers, reducing the model from 32 to 26 layers, and evaluated the models on the GSM8k dataset. For these experiments, we generated model responses using the baseline and pruned variants on the GSM8k dataset to capture how the distribution shift affects reasoning tasks post-pruning. Following Yang et al. (2024), to quantify the distribution shift, we employed Sentence-BERT (Reimers and Gurevych, 2019) to derive sentence embeddings from the model-generated responses. Then, similar to the method proposed by Zhang et al. (2023), we calculated the cosine similarity between the sentence embeddings of the pruned models and those generated by the original Llama3.1-8B Instruct model.
@@ D.1 Geometric Consistency and Application
Table 2: Model quality results for pruned Llama3.1-8B Instruct models across various pruning block sizes and fine-tuning strategies. This table reports the quality of different fine-tuning methods (No Fine-tuning, Standard Fine-tuning (SFT), and Self-Data Distillation) on various datasets, with average accuracy across ARC-C, GSM8k, and MMLU tasks. The "Avg. Recovery" column shows the percentage of model quality recovered relative to the unpruned baseline. The table highlights that the self-data distillation strategy consistently yields superior recovery rates, particularly with the 50k-sample OpenMathInstruct dataset. For instance, at a pruning block size of 6, the self-data distilled OpenMathInstruct model retains 95.96% of the original unpruned Llama3.1-8B Instruct (i.e., 32 layers) model’s quality, the highest recovery observed among all datasets and fine-tuning methods.

### v3

@@ 6 Conclusion
In conclusion, we introduce self-data distilled fine-tuning as an effective method to mitigate quality degradation in pruned LLMs, addressing catastrophic forgetting while preserving alignment with the model’s original data distribution. Our approach consistently outperforms standard supervised fine-tuning, demonstrating superior accuracy recovery post-pruning across various downstream tasks on the OpenLLM Leaderboard v1. Additionally, model merging via SLERP further enhances recovery, achieving significant quality retention. We also show that our method scales with dataset size, where larger self-distilled datasets lead to improved quality recovery. Moreover, integrating self-data distillation with speculative decoding not only enhances token acceptance rates but also reduces inference latency, offering an effective strategy for deploying pruned LLMs efficiently. These findings highlight self-data distilled fine-tuning as a critical tool for maintaining high model quality post-pruning, offering an efficient solution for model compression. Future work may involve integrating self-data distilled fine-tuning with complementary model compression techniques such as sparsity, quantization or teacher distillation, potentially yielding greater efficiency without sacrificing model quality. Extending these methodologies to next-generation LLM architectures presents a promising avenue for unlocking additional computational efficiency and model robustness.
@@ A.1 Baseline Model
In our ablation studies in Section 3, we used Llama3.1-8B Instruct [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct(Dubey et al., 2024) as the baseline model for all experiments. This model comprises a total of 32 decoder layers, pretrained on a diverse array of instruction-following datasets. This model was chosen for its strong generalization performance across a wide range of natural language processing (NLP) tasks, making it an ideal candidate for studying the impact of structured pruning and fine-tuning. The 8B model size strikes a balance between computational efficiency and model quality, providing a robust foundation for the experiments in this study. Hence, served as the starting point for our structured layer pruning ablations and experiments in Sections 3 and 4, respectively.
In Section 4, to further understand the efficacy of our methodology on other LLMs, we also applied it to Mistral-7B Instruct v0.3  [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, an instruct fine-tuned version of Mistral-7B-v0.3. This most recent version of Mistral-7B (Jiang et al., 2023), compared to Mistral-7B-v0.2, includes an extended vocabulary of 32,768 tokens, supports a v3 tokenizer, and enables function calling.
@@ A.2 Structured Layer Pruning
A.3 Calibration Dataset for Structured Layer Pruning
In structured pruning, selecting a suitable calibration dataset is critical for effectively identifying and removing redundant layers without sacrificing model quality. This study examines the impact of different calibration datasets: C4, RedPajama, and SlimPajama on computing the angular cosine distance block importance metric for Llama3.1-8B Instruct, as shown in Table 7. We found that all three datasets produced similar results across various prune block sizes, with comparable layers identified for removal. Given the consistency of results, we opted to use Redpajama as the calibration dataset in subsequent experiments due to its representative performance and alignment with our goals for efficient model pruning. Recent studies around the time of this submission have explored the nuanced role of calibration data in pruning large language models (Ji et al., 2024). While our analysis focuses on the practical selection of calibration datasets, a deeper investigation into calibration dataset characteristics and their influence on pruning decisions remains an open question for future work.
Table 7: Ablation study on the choice of calibration dataset for computing the angular cosine distance block importance metric on Llama3.1-8B Instruct. For calibration, we use a subset of 128 samples at a maximum sequence length (MSL) of 4096 from each dataset. The datasets C4, Redpajama, and Slimpajama produced similar pruning results across various block sizes, with comparable layers removed. Based on these results, the Redpajama dataset was selected for further evaluations due to its representative performance.
Block Size Removed Layers Dataset Score (avg dist) 2 24-25 C4 0.145 23-24 Redpajama 0.168 24-25 Slimpajama 0.153 4 24-27 C4 0.197 23-26 Redpajama 0.222 23-26 Slimpajama 0.205 6 22-27 C4 0.241 22-27 Redpajama 0.270 23-28 Slimpajama 0.249 8 20-27 C4 0.282 20-27 Redpajama 0.293 20-27 Slimpajama 0.289
A.4 Fine-tuning Datasets
@@
  • Alpaca Cleaned [^3] https://huggingface.co/datasets/yahma/alpaca-cleaned (Taori et al., 2023) The Alpaca Cleaned dataset is a cleaned version of the original Stanford Alpaca dataset, containing 51,760 instruction-following examples. It addresses several issues present in the original release, such as hallucinations, incorrect instructions, and output inconsistencies. This dataset provides high-quality general instruction-following tasks, spanning text generation, summarization, reasoning, and more. The cleaned version offers improved consistency and accuracy, making it ideal for fine-tuning large models in real-world instruction-following tasks.
  • OpenMathInstruct (Toshniwal et al., 2024) The OpenMathInstruct-1 dataset is specifically designed for fine-tuning language models on mathematical instruction tasks. It contains 1.8 million problem-solution pairs, generated using Mixtral-8x7B (Jiang et al., 2024). The problem sets are drawn from well-established mathematical benchmarks, including the GSM8K and MATH (Hendrycks et al., 2021b) datasets, ensuring a diverse and challenging range of mathematical reasoning tasks. Solutions are generated synthetically by allowing the Mixtral model to leverage a combination of natural language reasoning and executable Python code, which allows for both symbolic computation and procedural solutions. This combination of text and code execution makes the dataset particularly suited for training models to handle complex reasoning, problem-solving, and algebraic tasks.
A.4.1 Data Sampling and Experimental Consistency
@@
A.5 Fine-tuning Pruned Models
@@
Through this grid search, the optimal configuration was identified as a rank size = 8, epochs = 5, a batch size = 64, and learning rate = 1$\times$10-4. These hyperparameters were used consistently across all fine-tuning experiments (i.e., both standard supervised fine-tuning and self-data distilled fine-tuning) in this study to ensure a fair comparison of the models and their quality post-pruning. We conduct our model training using LLaMA-Factory v0.8.3444https://github.com/hiyouga/LLaMA-Factory/releases/tag/v0.8.3, a versatile framework designed for large-scale language model training and fine-tuning. This version offers extensive support for efficient parallelism, optimized memory usage, and integration with popular datasets, making it ideal for large model fine-tuning tasks such as those performed in this study.
A.5.1 Computational Resources
@@ Appendix C Experimental Setup for Understanding Catastrophic Forgetting
To understand the impact of distribution shift on catastrophic forgetting, we conducted experiments using the baseline model (i.e., Llama3.1-8B Instruct) and its pruned variants fine-tuned with both supervised fine-tuning (SFT) and self-data distilled fine-tuning (Self-Data FT). Specifically, we pruned 6 decoder layers, reducing the model from 32 to 26 layers, and evaluated the models on the GSM8k dataset. For these experiments, we generated model responses using the baseline and pruned variants on the GSM8k dataset to capture how the distribution shift affects reasoning tasks post-pruning. Following Yang et al. (2024), to quantify the distribution shift, we employed Sentence-BERT (Reimers & Gurevych, 2019) to derive sentence embeddings from the model-generated responses. Then, similar to the method proposed by Zhang et al. (2023), we calculated the cosine similarity between the sentence embeddings of the pruned models and those generated by the original Llama3.1-8B Instruct model.
@@ D.1 Geometric Consistency and Application
Table 8: Model quality results for pruned Llama3.1-8B Instruct models across various pruning block sizes and fine-tuning strategies. This table reports the quality of different fine-tuning methods (No Fine-tuning, Standard Fine-tuning (SFT), and Self-Data Distillation) on various datasets, with average accuracy across ARC-C, GSM8k, and MMLU tasks. The ”Avg. Recovery” column shows the percentage of model quality recovered relative to the unpruned baseline. The table highlights that the self-data distillation strategy consistently yields superior recovery rates, particularly with the 50k-sample OpenMathInstruct dataset. For instance, at a pruning block size of 6, the self-data distilled OpenMathInstruct model retains 95.96% of the original unpruned Llama3.1-8B Instruct (i.e., 32 layers) model’s quality, the highest recovery observed among all datasets and fine-tuning methods.